Using the table below, show and explain how Mexico has the comparative advantage in phones.

  

Mexico

United States

Possibility

Burgers

Phones

Burgers

Phones

A

6

0

12

0

B

4

8

8

8

C

2

16

4

16

D

0

24

0

24
Answer #1

Solution -

we know that ,

comparative advantage means compare advantage between to points

Now Find For Mexico,

here we put inputs

6/6 = 24/6

1 burger = 4 phones

Now Find For United States,

here we put inputs

= 12/12 = 24/121

1 burger = 2 phones